OpenStudy (anonymous):

what is the perimeter of a rectangle with a length of 4 1/4 inches and the width of 2 5/6 A. 6 3/5 B. 7 1/12 C. 13 1/5 D. 14 1/6

OpenStudy (anonymous):

will give medals will fan

OpenStudy (dtan5457):

double the length and double the width. add the two together.

OpenStudy (dtan5457):

\[4\frac{ 1 }{ 4 }=\frac{ 17 }{ 4 }\] \[\frac{ 17 }{ 4 }+\frac{ 17 }{ 4 }=?\]

OpenStudy (dtan5457):

\[2\frac{ 5 }{ 6 }=\frac{ 17 }{ 6 }\] \[\frac{ 17 }{ 6 }+\frac{ 17 }{ 6 }=?\]

OpenStudy (dtan5457):

Once you get those values, add those 2 fractions together.

OpenStudy (dtan5457):

and you get your answer :)